# Higher-weight Jacobians

### v4

@@ 1. Introduction
We see that a pure integral Hodge structure $H$ of weight $m>0$ has a Jacobian if $\mathfrak{L}(H)$ is a discrete subgroup of the complex vector space $H^{0,m}$. If $H={\rm H}^{m}(X,\mathbf{Z})$ with $X$ a complex variety, we instead say that $X$ has a Jacobian of weight $m$, or an $m$-Jacobian. In this case, $\mathfrak{C}_{m}(X)$ is naturally a complex torus. When $m=2$, we will also use the name Brauer-Jacobian, and we will write $\mathfrak{C}$ for $\mathfrak{C}_{2}$.
In particular, when $C$ is a curve, the “weight $1$ Jacobian of $C$” defined above coincides with the realization of the Jacobian of $C$ as a complex torus via the exponential sequence. Furthermore, note that there is a natural inclusion
$\ker({\rm H}^{m}(X,\mathbf{G}_{m})\to{\rm H}^{m+1}(X,\mathbf{Z}(1)))\hookrightarrow\mathfrak{C}_{m}(X)$
that identifies the former with the torsion subgroup of the latter. That is, the topologically trivial degree $m$ cohomology is identified with the torsion points of the $m$-Jacobian; in particular, when $X$ has a $2$-Jacobian or Brauer-Jacobian, the torsion points of the Brauer-Jacobian are identified with the Brauer group of $X$ (whence the name).
In this paper, we study examples of Brauer-Jacobians for several types of varieties, with a focus on CM abelian varieties, where we can make the calculations quite explicit. This general phenomenon has been studied before, notably in a paper of Beauville [beauville], where various surfaces of maximal Picard number were discussed. There is also the fascinating work of Kollár and Joglar-Prieto [kollar2003] in the context of real algebraic geometry.
@@
theoremkummercon If $\mathcal{A}$ is a complex abelian variety with an $m$-Jacobian $\mathfrak{C}_{m}(\mathcal{A})$, then the Kummer variety $\mathcal{K}(\mathcal{A})$ associated to $\mathcal{A}$ also has an $m$-Jacobian admitting a canonical isomorphism $\mathfrak{C}_{m}(\mathcal{K}(\mathcal{A}))\cong\mathfrak{C}_{m}(\mathcal{A})$.
@@
$[\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(\mathcal{O}),$
@@
For such varieties $X$, we end up with a complex torus $B$ and a “natural” inclusion $\operatorname{Br}(X)\hookrightarrow B$ as the torsion points. If $B$ is itself algebraic and sufficiently functorial, then one could use this geometric $2$-Jacobian for $\operatorname{Br}(X)$ to study it using the geometric and arithmetic properties of rational points on varieties. This would have many significant consequences, including applications to the Tate conjecture.
@@
propositiondecompfieldofdef Assume $\mathcal{A}=\mathbf{C}^{2}/\Lambda$ is a complex torus of dimension $2$ with a Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ such that $\mathcal{A}$ is primitive. Then there exist elliptic curves $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ such that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if the group element $[\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O}(\mathcal{A}))$ has order $\leq 2$.
Given that the construction of $\mathfrak{C}_{m}(X)$ appears to be “morally wrong”, it is somewhat bizarre that we keep finding algebraic tori (both here and in [beauville]). Note that there are some cases described in Remark 7.6 of [shioda72] where it is not clear if the torus is algebraic. One is naturally led to the following question.
@@ 2. Generalities on higher-weight Jacobians
In this section, we briefly discuss higher-weight Jacobians in general, then discuss applications for blowups and equivariant $m$-Jacobians. Fix a smooth, proper, complex variety $X$.
@@
Let $V$ be a pure real Hodge structure of weight $m$ with Hodge decomposition $V_{\mathbf{C}}=\bigoplus V^{i,j}$. For any $i$ and $j$ such that $i+j=m$, the $\mathbf{R}$-linear map $V\to V^{i,j}$ (from a real vector space to a complex vector space) is surjective.
@@
By construction we have $\mathfrak{C}(H)=\bigoplus\mathfrak{C}(H_{i})$, so the result follows immediately from Lemma 2.5. ∎
@@
Let $E$ be the preimage of $Y$ in $\widetilde{X}$. By Theorem 7.31 of [voisinthing], there is an isomorphism of Hodge structures
@@
To see the final assertion, note that (since smooth varieties have rational singularities) the natural map ${\rm H}^{m}(X,\mathscr{O}_{X})\to{\rm H}^{m}(\widetilde{X},\mathscr{O}_{\widetilde{X}})$ is an isomorphism that induces a map of $m$-Jacobians $\mathfrak{C}_{m}(X)\to\mathfrak{C}_{m}(\widetilde{X})$. The result follows. ∎
@@
Given two birational varieties $X$ and $X^{\prime}$ of dimension at most $4$, we have that $X$ has a $q$-Jacobian if and only if $X^{\prime}$ has a $q$-Jacobian, for any $q$.
@@
We can compare $X$ and $X^{\prime}$ by birational morphisms having centers of dimension at most $2$. Following the chain of morphisms and using Proposition 2.8 gives the desired result. ∎
@@
If $H$ has a Jacobian then the invariant substructure $H^{G}$ has a Jacobian.
@@
By flatness, we have $(H^{G})_{\mathbf{Q}}=(H_{\mathbf{Q}})^{G}$. Thus, it suffices to consider the rational Hodge structure $H_{\mathbf{Q}}$. By semisimplicity, we have a decomposition $H_{\mathbf{Q}}=H_{\mathbf{Q}}^{G}\oplus H^{\prime}$ of Hodge structures. The result now follows from Corollary 2.7. ∎
@@
Let $X$ be a smooth projective variety with an action by a finite group $G$ such that the quotient $Y=X/G$ is smooth. If $X$ has a $q$-Jacobian then $Y$ has a $q$-Jacobian.
@@
By the Cartan-Leray spectral sequence (Theorem XVI.8.4 of [cartaneilenberg]), we have
@@
Suppose $Y\dashrightarrow X$ is a rational double cover of smooth algebraic surfaces. Then for any $m$, $X$ has an $m$-Jacobian if and only if $Y$ has an $m$-Jacobian.
@@
By Corollary 2.9, Proposition 2.11, and the fact that any quadratic field extension (in characteristic different from $2$) is a $\mathbf{Z}/2\mathbf{Z}$-Galois extension, we see that is suffices to show that there is a birational model $\widetilde{X}\to X$ such that the pullback of $Y$ to $\widetilde{X}$ extends to a double cover $\widetilde{Y}\to\widetilde{X}$ with $\widetilde{Y}$ also smooth. To do this, note that we can make a birational modification $X^{\prime}\to X$ so that the normalization of $X^{\prime}$ is the function field of $Y$ gives a finite morphism $Y^{\prime}\to X^{\prime}$ with branch locus $B\subset X^{\prime}$ an snc divisor. If $B$ is smooth then so is $Y^{\prime}$. Otherwise, $B$ has simple crossings of curves, so $Y^{\prime}$ has $A_{1}$-singularities over the singular points of $B$. Blowing up these points in $X^{\prime}$ to yield $\widetilde{X}\to X^{\prime}$, we can explicitly calculate that the normalization $\widetilde{Y}\to\widetilde{X}$ is now smooth. ∎
We end this section with a comment on the relation between higher-weight Jacobians and the framework of [totaro].
@@ 3. Complex lattices and complex tori
Lattices in the complex numbers are closely related to complex elliptic curves; this is discussed in various sources including [silverman09]. In particular, by [silverman09, Corollary VI.5.1.1], every complex elliptic curve $E$ is isomorphic as a complex Lie group to $\mathbf{C}/\Lambda$ for some lattice $\Lambda\subset\mathbf{C}$, which is unique up to homothety.
Furthermore, by [silverman09, Theorem VI.4.1], every isogeny of complex elliptic curves $\phi:E_{1}\to E_{2}$ (with $E_{i}=\mathbf{C}/\Lambda_{i}$) arises from a homothety of $\Lambda_{1}$ with a sublattice of $\Lambda_{2}$—i.e., the isogeny corresponds to a nonzero complex number $\alpha$ with $\alpha\Lambda_{1}\subseteq\Lambda_{2}$. We will refer to lattices with this property as isogenous, as we do for the corresponding elliptic curves. Isogeny of lattices is an equivalence relation.
@@ 3.2. Field of definition
There exists a correspondence between complex elliptic curves and complex lattices, as discussed in previous sections. We note that two elliptic curves defined over number fields are isomorphic over $\mathbf{C}$ if and only if they have the same $j$-invariant; we generally treat the $j$-invariant of an elliptic curve and the $j$-invariant of the associated lattice as interchangeable [silvermanAd, Proposition 4.4].
@@
and it is easy to find examples of elliptic curves over $\mathbf{Q}$ with $j$-invariant 00 or $1728$.
@@
We will give a criterion in Proposition 1.2 for when a complex abelian surface $\mathcal{A}$ having a Brauer-Jacobian $\mathfrak{C}(A)$ can be defined over the same field as the elliptic curve $\mathfrak{C}(A)$.
The first main theorem of complex multiplication [cox11, Theorem 11.1] states that for an order $\mathcal{O}$ of an imaginary quadratic field $K$, the $j$-invariant of any proper fractional $\mathcal{O}$-ideal $I$ (or equivalently, the $j$-invariant of any lattice with CM by $\mathcal{O}$) is an algebraic integer $j(I)$, and for all $I$ the field extension $K(j(I))$ is the ring class field of the order $\mathcal{O}$.
The ring class field $L_{\mathcal{O}}$ has the property that $\operatorname{Cl}(\mathcal{O})\simeq{\rm Gal}(L_{\mathcal{O}}/K)$. Under this group isomorphism, the class of a proper fractional ideal $I$ corresponds to the element $\sigma_{I}$ of the Galois group which permutes $j$-invariants by $[\overline{I}]=[I^{-1}]$—i.e., $\sigma_{I}(j(J))=j(I^{-1}J)=j(\overline{I}J))$ for $I,J$ proper fractional ideals of $\mathcal{O}$ [cox11, Corollary 11.37].
@@
    or $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$ is a degree $n$ extension of $\mathbf{Q}$, and in $\operatorname{Cl}(\mathcal{O})$ the element $[E_{1}][E_{2}]^{-1}$ has order $2$.
@@
Writing $\mathbf{Z}/2\mathbf{Z}$ as {±$\{\pm 1\}$, the order $2$ elements of $\operatorname{Cl}(\mathcal{O})\rtimes\mathbf{Z}/2\mathbf{Z}$ are either $(x,1)$ for $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$, or $([E],-1)$ for any $[E]\in\operatorname{Cl}(\mathcal{O})$. Since $\operatorname{Cl}(\mathcal{O})$ acts transitively and freely on the set of $j$-invariants of elliptic curves with CM by $\mathcal{O}$ [cox11, Corollary 11.37], the subfield $\mathbf{Q}(j(E_{1}))$ cannot arise as the fixed field of $(x,1)$ for any $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$.
Therefore for any elliptic curve $E$ with ${\rm End}(E)=\mathcal{O}$, the subextension $\mathbf{Q}(j(E))\subseteq L_{\mathcal{O}}$ is the fixed field of $([I],-1)$ for some proper fractional ideal $I$ of $\mathcal{O}$. Using the description of the action of $\operatorname{Cl}(\mathcal{O})$ on the $j$-invariants of elliptic curves with CM by $\mathcal{O}$ given in [cox11, Corollary 11.37], we see that for any proper fractional ideal $J$ of $\mathcal{O}$, the element $([I],-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$ sends $j(J)$ to $j(\overline{IJ})=j(I^{-1}J^{-1})$. Therefore $([I],-1)$ fixes $j(J)$ if and only if $[J]^{2}=[I]^{-1}\in\operatorname{Cl}(\mathcal{O})$.
@@
$[\mathbf{Q}(j(E_{1}),j(E_{2})):\mathbf{Q}]=n\text{ or }2n.$
In the latter case, (i) obviously holds; in the former case, $\mathbf{Q}(j(E_{1}),j(E_{2}))=\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$, so $([E_{1}]^{-2},-1)=([E_{2}]^{-2},-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$. Then $[E_{1}]^{2}=[E_{2}]^{2}\in\operatorname{Cl}(\mathcal{O})$, meaning $[E_{1}][E_{2}]^{-1}$ has order $2$ (since $E_{1}\not\simeq E_{2}$) and we have (ii) in this case, completing the proof. ∎
@@
[…]
@@
[…]

### v5

@@ 1. Introduction
We see that a pure integral Hodge structure $H$ of weight $m>0$ has a Jacobian if $\mathfrak{L}(H)$ is a discrete subgroup of the complex vector space $H^{0,m}$. This gives $\mathfrak{C}(H)=H^{0,m}/\mathfrak{L}(H)$ a natural structure of a compact complex torus $\mathbf{C}^{g}/\Gamma$. [^1] Throughout this paper, we use “complex torus” to refer to a compact complex torus. If $H={\rm H}^{m}(X,\mathbf{Z})$ with $X$ a complex variety, we instead say that $X$ has a Jacobian of weight $m$, or an $m$-Jacobian. When $m=2$, we will also use the name Brauer-Jacobian, and we will write $\mathfrak{C}$ for $\mathfrak{C}_{2}$.
In particular, when $C$ is a curve, the “weight $1$ Jacobian of CC” defined above coincides with the realization of the Jacobian of $C$ as a complex torus via the exponential sequence. Furthermore, note that there is a natural map
[…]
In this paper, we study examples of Brauer-Jacobians for several types of varieties, with a focus on CM abelian varieties, where we can make the calculations quite explicit. This general phenomenon has been studied before, notably in a paper of Beauville [Bea14], where various surfaces of maximal Picard number were discussed. There is also the fascinating work of Kollár and Joglar-Prieto [JK03] in the context of real algebraic geometry.
@@
theoremkummercon If $\mathcal{A}$ is a complex abelian variety with an $2m$-Jacobian $\mathfrak{C}_{2m}(\mathcal{A})$, then the Kummer variety $\mathcal{K}(\mathcal{A})$ associated to $\mathcal{A}$ also has an $2m$-Jacobian admitting a canonical isomorphism $\mathfrak{C}_{2m}(\mathcal{K}(\mathcal{A}))\cong\mathfrak{C}_{2m}(\mathcal{A})$.
@@
$\mathfrak{C}(E_{1}\times E_{2})]=[E_{1}][E_{2}]\in\operatorname{Cl}(\mathcal{O}),$
@@
For such varieties $X$, we end up with a complex torus $B$ and a “natural” map from $\operatorname{Br}(X)$ to the torsion points of $B$. If $B$ is itself algebraic and sufficiently functorial, then one could use this geometric $2$-Jacobian for $\operatorname{Br}(X)$ to study it using the geometric and arithmetic properties of rational points on varieties. This would have many significant consequences, including applications to the Tate conjecture.
@@
propositiondecompfieldofdef Assume $\mathcal{A}=\mathbf{C}^{2}/\Lambda$ is a complex torus of dimension $2$ with a Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$ such that $\mathcal{A}$ is primitive. Then there exist elliptic curves $E_{1},E_{2}$ with $\mathcal{A}\simeq E_{1}\times E_{2}$ such that $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))=\mathbf{Q}(j(\mathfrak{C}(\mathcal{A})))$ if and only if the group element $\mathfrak{C}(\mathcal{A})]\in\operatorname{Cl}(\mathcal{O}(\mathcal{A}))$ has order $\leq 2$.
Given that the construction of $\mathfrak{C}_{m}(X)$ appears to be “morally wrong”, it is somewhat bizarre that we keep finding algebraic tori (both here and in [Bea14]). Note that there are some cases described in Remark 7.6 of [Shi72] where it is not clear if the torus is algebraic. One is naturally led to the following question.
@@ 2. Generalities on higher-weight Jacobians
In this section, we briefly discuss higher-weight Jacobians in general, then discuss applications for blowups and equivariant $m$-Jacobians. Fix a smooth, proper, complex variety $X$. We assume $m>0$ throughout this section.
@@
Let $V$ be a pure real Hodge structure of weight $m$ with Hodge decomposition $V_{\mathbf{C}}=\bigoplus V^{i,j}$. For any $i$ and $j$ such that $i\neq j,i+j=m$, the $\mathbf{R}$-linear map $V\to V^{i,j}$ (from a real vector space to a complex vector space) is surjective.
@@
By construction we have $\mathfrak{C}(H)=\bigoplus\mathfrak{C}(H_{i})$, so the result follows immediately from Remark 2.5. ∎
@@
Let $E$ be the preimage of $Y$ in $\widetilde{X}$. By Theorem 7.31 of [Voi02], there is an isomorphism of Hodge structures
@@
[…]
@@
Let $X$ and $X^{\prime}$ be birational smooth projective varieties. Then $X$ has a $q$-Jacobian if and only if $X^{\prime}$ has a $q$-Jacobian, for any $q$.
@@
By the Weak Factorization Theorem, we can compare $X$ and $X^{\prime}$ by a chain of blowups and blowdowns with smooth centers. Following the diagram of these morphisms and using Proposition 2.9 gives the desired result. ∎
@@
If $H$ has a Jacobian then the invariant substructure $H^{G}$ has a Jacobian; moreover, $\mathfrak{C}(H)$ admits a $G$-action and $\mathfrak{C}(H)^{G}=\mathfrak{C}(H^{G})$.
@@
By flatness, we have $(H^{G})_{\mathbf{Q}}=(H_{\mathbf{Q}})^{G}$. Thus, it suffices to consider the rational Hodge structure $H_{\mathbf{Q}}$. By semisimplicity, we have a decomposition $H_{\mathbf{Q}}=H_{\mathbf{Q}}^{G}\oplus H^{\prime}$ of Hodge structures.
The result now follows from Corollary 2.7. The Jacobian $\mathfrak{C}(H)$ naturally inherits the $G$-action of $H$, and by construction we see that $\mathfrak{C}(H)=\mathfrak{C}(H^{G})\oplus\mathfrak{C}(H^{\prime})$, with the former summand evidently the invariant part of $\mathfrak{C}(H)$. ∎
@@
Let $X$ be a smooth projective variety with an action by a finite group $G$ such that the quotient $Y=X/G$ is smooth. If $X$ has a $q$-Jacobian then $Y$ has a $q$-Jacobian. Morever, $\mathfrak{C}(X)$ admits a $G$-action and $\mathfrak{C}(X)^{G}=\mathfrak{C}(Y)$.
@@
By the Cartan-Leray spectral sequence (Theorem XVI.8.4 of [CE56]), we have
@@
Suppose $Y\dashrightarrow X$ is a rational double cover of smooth algebraic surfaces. Then for any $m$, $X$ has an $m$-Jacobian if $Y$ has an $m$-Jacobian.
@@
It suffices to show that there is a birational model $\widetilde{X}\to X$ such that the pullback of $Y$ to $\widetilde{X}$ extends to a double cover $\widetilde{Y}\to\widetilde{X}$ with $\widetilde{Y}$ also smooth. This is because $\widetilde{Y}$ and $Y$ will have isomorphic $m$-Jacobians by Corollary 2.10. Moreover, the double cover $\widetilde{Y}\to\widetilde{X}$ is a $\mathbf{Z}/2\mathbf{Z}$-Galois cover since any quadratic field extension (in characteristic different from $2$) is a $\mathbf{Z}/2\mathbf{Z}$-Galois extension, so we may apply Proposition 2.12 to see that $\widetilde{X}$ has an $m$-Jacobian. Finally we apply Corollary 2.10 again to see that $X$ has an $m$-Jacobian as we deisred to show.
To obtain the birational model $\widetilde{X}\to X$, note that we can make a birational modification $X^{\prime}\to X$ so that the normalization of $X^{\prime}$ in the function field of $Y$ gives a finite morphism $Y^{\prime}\to X^{\prime}$ such that its branch locus $B\subset X^{\prime}$ is an snc divisor. If $B$ is smooth then so is $Y^{\prime}$.
Otherwise, $B$ has simple crossings of curves, so $Y^{\prime}$ has $A_{1}$-singularities over the singular points of $B$. Blowing up these points in $X^{\prime}$ to yield $\widetilde{X}\to X^{\prime}$, we can explicitly calculate that the normalization $\widetilde{Y}\to\widetilde{X}$ is now smooth. ∎
We end this section with a comment on the relation between higher-weight Jacobians and the framework of [Tot15].
@@ 3. Complex lattices and complex tori
Lattices in the complex numbers are closely related to complex elliptic curves; this is discussed in various sources including [Sil09]. In particular, by [Sil09, Corollary VI.5.1.1], every complex elliptic curve $E$ is isomorphic as a complex Lie group to $\mathbf{C}/\Lambda$ for some lattice $\Lambda\subset\mathbf{C}$, which is unique up to homothety.
Furthermore, by [Sil09, Theorem VI.4.1], every isogeny of complex elliptic curves $\phi:E_{1}\to E_{2}$ (with $E_{i}=\mathbf{C}/\Lambda_{i}$) arises from a homothety of $\Lambda_{1}$ with a sublattice of $\Lambda_{2}$—i.e., the isogeny corresponds to a nonzero complex number $\alpha$ with $\alpha\Lambda_{1}\subseteq\Lambda_{2}$. We will refer to lattices with this property as isogenous, as we do for the corresponding elliptic curves. Isogeny of lattices is an equivalence relation.
@@ 3.2. Field of definition
There exists a correspondence between complex elliptic curves and complex lattices, as discussed in previous sections. We note that two elliptic curves defined over number fields are isomorphic over $\mathbf{C}$ if and only if they have the same $j$-invariant; we generally treat the $j$-invariant of an elliptic curve and the $j$-invariant of the associated lattice as interchangeable [Sil94, Proposition 4.4].
@@
and it is easy to find examples of elliptic curves over $\mathbf{Q}$ with $j$-invariant 0 or $1728$.
@@
We will use the results of this section to relate the field of definition of a complex abelian surface $\mathcal{A}$ to that of its Brauer-Jacobian $\mathfrak{C}(\mathcal{A})$, which is an elliptic curve (Proposition 1.2).
The first main theorem of complex multiplication [Cox11, Theorem 11.1] states that for an order $\mathcal{O}$ of an imaginary quadratic field $K$, the $j$-invariant of any proper fractional $\mathcal{O}$-ideal $I$ (or equivalently, the $j$-invariant of any lattice with CM by $\mathcal{O}$) is an algebraic integer $j(I)$, and for all proper ideals $I\subseteq\mathcal{O}$ the field extension $K(j(I))$ is the ring class field of the order $\mathcal{O}$.
The ring class field $L_{\mathcal{O}}$ has the property that $\operatorname{Cl}(\mathcal{O})\simeq{\rm Gal}(L_{\mathcal{O}}/K)$. Under this group isomorphism, the class of a proper fractional ideal $I$ corresponds to the element $\sigma_{I}$ of the Galois group which permutes $j$-invariants by $\overline{I}]=[I^{-1}]$—i.e., $\sigma_{I}(j(J))=j(I^{-1}J)=j(\overline{I}J))$ for $I,J$ proper fractional ideals of $\mathcal{O}$ [Cox11, Corollary 11.37].
@@
    or $\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$ is a degree $n$ extension of $\mathbf{Q}$, and in $\operatorname{Cl}(\mathcal{O})$ the element $E_{1}][E_{2}]^{-1}$ has order $2$.
@@
Writing $\mathbf{Z}/2\mathbf{Z}$ as $\{\pm 1\}$, the order $2$ elements of $\operatorname{Cl}(\mathcal{O})\rtimes\mathbf{Z}/2\mathbf{Z}$ are either $(x,1)$ for $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$, or ([E],1)([E],-1) for any $E]\in\operatorname{Cl}(\mathcal{O})$. Since $\operatorname{Cl}(\mathcal{O})$ acts transitively and freely on the set of $j$-invariants of elliptic curves with CM by $\mathcal{O}$ [Cox11, Corollary 11.37], the subfield $\mathbf{Q}(j(E_{1}))$ cannot arise as the fixed field of $(x,1)$ for any $x\in\operatorname{Cl}(\mathcal{O})$ of order $2$.
Therefore for any elliptic curve $E$ with ${\rm End}(E)=\mathcal{O}$, the subextension $\mathbf{Q}(j(E))\subseteq L_{\mathcal{O}}$ is the fixed field of ([I],1)([I],-1) for some proper fractional ideal $I$ of $\mathcal{O}$. Using the description of the action of $\operatorname{Cl}(\mathcal{O})$ on the $j$-invariants of elliptic curves with CM by $\mathcal{O}$ given in [Cox11, Corollary 11.37], we see that for any proper fractional ideal $J$ of $\mathcal{O}$, the element $([I],-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$ sends $j(J)$ to $j(\overline{IJ})=j(I^{-1}J^{-1})$. Therefore ([I],1)([I],-1) fixes $j(J)$ if and only if $J]^{2}=[I]^{-1}\in\operatorname{Cl}(\mathcal{O})$.
@@
[…]
In the latter case, (i) obviously holds; in the former case, $\mathbf{Q}(j(E_{1}),j(E_{2}))=\mathbf{Q}(j(E_{1}))=\mathbf{Q}(j(E_{2}))$, so $([E_{1}]^{-2},-1)=([E_{2}]^{-2},-1)\in{\rm Gal}(L_{\mathcal{O}}/\mathbf{Q})$. Then $E_{1}]^{2}=[E_{2}]^{2}\in\operatorname{Cl}(\mathcal{O})$, meaning $E_{1}][E_{2}]^{-1}$ has order $2$ (since $E_{1}\not\simeq E_{2}$) and we have (ii) in this case, completing the proof. ∎
@@
The order $\mathfrak{O}=\mathbf{Z}[\sqrt{-27}]$ in the field $\mathfrak{K}=\mathbf{Q}(\sqrt{-3})=\mathbf{Q}(\zeta_{3})$ (where $\zeta_{3}=\frac{-1+\sqrt{-3}}{2}$ is a cube root of unity) has ring class field $L_{\mathfrak{O}}=\mathfrak{K}(\sqrt[3]{2})=\mathbf{Q}(\zeta_{3},\sqrt[3]{2})$ and class group $\operatorname{Cl}(\mathfrak{O})=\mathbf{Z}/3\mathbf{Z}$. The three homothety classes of lattices with CM by $\mathfrak{O}$ can be computed to be represented by
@@
The order $\mathfrak{R}^{\prime}=\mathbf{Z}[\sqrt{-36}]$ in the field $K=\mathbf{Q}(i)$ has ring class field $L_{\mathfrak{R}}^{\prime}=K(\sqrt[4]{12})$ and class group $\operatorname{Cl}(\mathfrak{R}^{\prime})=\mathbf{Z}/4\mathbf{Z}$. The four homothety classes of lattices with CM by $\mathfrak{R}^{\prime}$ can be computed to be represented by $\mathfrak{R}^{\prime}=\langle 1,6i\rangle,\langle 3,2i\rangle,\langle 3,1\pm 2i\rangle$ (Example A.2). The lattices $\langle 1,6i\rangle$ and $\langle 3,2i\rangle$ both have real $j$-invariants generating the extension $\mathbf{Q}(\sqrt[4]{12})/\mathbf{Q}$. The lattices $\langle 3,1\pm 2i\rangle$ both have $j$-invariants generating the extension $\mathbf{Q}(i\sqrt[4]{12})/\mathbf{Q}$.